OpenStudy (anonymous):

∫5sec (π/4−x) tan (π/4−x) dx....does anyone know how to solve this....?

OpenStudy (anonymous):

\[put \frac{ \pi }{ 4 }-x=t,-dx=dt\] \[\int\limits \sec x \tan x dx=\sec x+c\]

OpenStudy (anonymous):

can u explain more?

OpenStudy (anonymous):

\[I=5\int\limits \sec \left( \frac{ \pi }{4 }-x \right)\tan \left( \frac{ \pi }{4 }-x \right)dx\] \[put \frac{ \pi }{ 4 }-x=t,-dx=dt,dx=-dt\] \[I=-5\int\limits \sec t \tan t dt=-5\sec t+c\] \[=-5\sec \left( \frac{ \pi }{ 4}-x \right)+c\]